OpenStudy (anonymous):

4 (before square root ) square root 256x^6

OpenStudy (tkhunny):

\(4\sqrt{256x^{6}}\) Find perfect squares inside the radical. \(256 = 16^{2}\) \(x^{2} = \left(x^{3}\right)^{2}\) Simplify away...

OpenStudy (anonymous):

\[\sqrt[4]{256x^6?}\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Why is that different? Find perfect 4ths: \(256 = 4^{4}\) \(x^{6} = x^{4}\cdot x^{2}\) <== Be VERY careful with this piece.

OpenStudy (anonymous):

sorry i wrote it wrong the first time

OpenStudy (tkhunny):

Okay, have at it.

OpenStudy (anonymous):

\[\sqrt[4]{4x^6?}\]

OpenStudy (anonymous):

since 256 =4^4 the 4 before the square root will cancel out

OpenStudy (anonymous):

right? @tkhunny

OpenStudy (anonymous):

wait would it be 1/4x^1/2 or 4/x^1/2

OpenStudy (anonymous):

someone please help

OpenStudy (tkhunny):

?? \(\sqrt[4]{256} = \sqrt[4]{4^{4}} = 4^{4/4} = 4^{1} = 4\) You do the 'x' part.

OpenStudy (anonymous):

x^1/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

or would it be 4x^3/2?

OpenStudy (tkhunny):

Why not follow the nice, clear path? \(\sqrt[4]{x^{6}} = \sqrt[4]{x^{4}\cdot x^{2}} = \sqrt[4]{x^{4}}\cdot\sqrt[4]{x^{2}} = x^{4/4}\cdot x^{2/4} = |x|\cdot\sqrt{|x|}\) It's a little tricky towards the end.

OpenStudy (anonymous):

x multiply by x^1/2

OpenStudy (tkhunny):

Well, the absolute values are necessary, but your particular exam or teacher may not care about that.

OpenStudy (anonymous):

which equals x^3/2

OpenStudy (anonymous):

4x^3/2 would be the answer

OpenStudy (tkhunny):

That is one possible correct form.

OpenStudy (anonymous):

so it is correct?